OpenStudy (anonymous):

Solve 7.7^4x + 3 = 10.2 x ≈ –2.15 x ≈ –0.47 x ≈ 0.03 x ≈ 0.28

OpenStudy (nottim):

ER.....Isolate x? Use logs. u kno wat logs are?

OpenStudy (anonymous):

I just need this question done to prove that my mom is wrong so I want to see what every one gets as an answer. I don't know if I did it correctly though..

OpenStudy (nottim):

well, show yur steps and answer. ill get some others to come check it out once you've posted it.

OpenStudy (anonymous):

7.7^(4x+3) = 10.2 (4x+3) log(7.7) = log(10.2) 4x + 3 = log(10.2) / log(7.7) 4x = log(10.2) / log(7.7) − 3 x = log(10.2) / (4 log(7.7)) − 3/4 x = −0.465563812 but it doesn't match any of the answers.. unless -.047 would be it?

OpenStudy (lgbasallote):

seems right to me